What is the initial velocity of an automobile that accelerate at 5.0m/s² and has displacement of 114m in 6.0s?


1
Expert's answer
2021-07-27T16:46:03-0400
"s=vt+0.5at^2\\114=6v+0.5(5)(6)^2\\v=4\frac{m}{s}"